[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Wed Sep 21, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Thu Sep 22 02:23:05 UTC 2016


HSU Shift Log for Wed Sep 21 2016 22:22:58 GMT-0400 (EDT)

Shift: R

Forecaster: D Brown

Technical Issues Summary:

ATCF issues:

N-AWIPS/AWIPS2 issues: Hi-resolution ASCAT-A and ambiguities have not  
arrived since about 21z. Low resolution data available and ASCAT B data  
coming in. This is similar to what occurred on Monday.

Other Techincal issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160922/0d1f28f3/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list